# Covering rings by proper ideals

### v1

@@ 4. An infinite family of $\eta_{\ell}$-elementary rings
For each $v\in\mathbb{F}_{q}^{n}$ (viewed as a column vector) and linear subspace $V\subseteq\mathbb{F}_{q}^{n}$, let $L_{v}:=\{(A|Av):A\in S\}$ and $N_{V}:=\{(A|w):AV=0\text{ and }w\in\mathbb{F}_{q}^{n}\}$. That is, $N_{V}$ consists of the elements $(A|w)\in R$ such that $V$ is contained in the right nullspace of $A$, and there are no restrictions on $w$.
@@
    For each linear subspace $V\subseteq\mathbb{F}_{q}^{n}$, $N_{V}$ is a maximal left ideal of $R$.
@@
    For each linear subspace $V\subseteq\mathbb{F}_{q}^{n}$, $N_{V}$ is a cyclic left ideal of $R$.
@@
    Let $V$ and $W$ be linear subspaces of $\mathbb{F}_{q}^{n}$. Then, $N_{V}=N_{W}$ if and only if $V=W$.
@@
    Let $(A|v)\in R$ such that $A\in S$ is not invertible. Then, $(A|v)\in N_{V}$ for any linear subspace $V\subseteq\mathbb{F}_{q}^{n}$ such that $AV=0$.
@@
(1) Let $V$ be a linear subspace of $\mathbb{F}_{q}^{n}$. It is easy to verify that $N_{V}$ is a left ideal of $R$. For maximality, let $V^{\perp}$ be the subspace complement to $V$ in $\mathbb{F}_{q}^{n}$. Then, $\dim_{\mathbb{F}_{q}}(V^{\perp})=n-1$, so $|V^{\perp}|=q^{n-1}$. Now, $(A|w)\in N_{V}$ if and only if each row of $A$ is the transpose of a vector in $V^{\perp}$. Thus, there are $q^{(n-1)n}$ choices for $A$ and $q^{n}$ choices for $w$, so $|N_{V}|=q^{(n-1)n}q^{n}=q^{n^{2}}$. Thus, $N_{V}$ is maximal by Lemma 4.2.
(2) Fix a linear subspace $V\subseteq\mathbb{F}_{q}^{n}$ with complement subspace $V^{\perp}$. Let $\{b_{1},\ldots,b_{n-1}\}$ be a basis (of column vectors) for $V^{\perp}$. Let
@@
Then, $b_{i}^{T}v=0$ for all $1\leqslant i\leqslant n-1$ and all $v\in V$, so $B\in N_{V}$. Next, given, $(A|w)\in N_{V}$, the transpose of row $i$ of $A$ is in $V^{\perp}$, so that row is an $\mathbb{F}_{q}$-linear combination of $b_{1}^{T},\ldots,b_{n-1}^{T}$. For $1\leqslant i\leqslant n$ and $1\leqslant j\leqslant n-1$, let $c_{i,j}\in\mathbb{F}_{q}$ be such that row $i$ of $A$ is equal $c_{i,1}b_{1}^{T}+\cdots+c_{i,n-1}b_{n-1}^{T}$. Also, let $w_{i}$ be the entry of $w$ in row $i$. Then,
@@
(3) This follows from the observation that for any linear subspaces $V$ and $W$, there exist matrices $A_{V}$ and $A_{W}$ such that the right nullspaces of $A_{V}$ and $A_{W}$ are exactly $V$ and $W$, respectively.
@@
By Lemmas 4.4(1), 4.5(1) and 4.5(2), each left ideal $L_{v}$ and $N_{V}$ is a maximal left ideal of $R$, and is generated (as a left ideal) by a single element. Thus, each $L_{v}$ and $N_{V}$ must be part of any cover of $R$ by left ideals. By Lemma 4.4(2), the number of distinct maximal left ideals $L_{v}$ is $|\mathbb{F}_{q}^{n}|=q^{n}$; and by Lemma 4.5(3), the number of distinct maximal left ideals $N_{V}$ is equal to the number of linear subspaces of $\mathbb{F}_{q}^{n}$, which is (using Gaussian binomial coefficients) $\binom{n}{1}_{q}=1+q+\cdots+q^{n-1}$. So, $\eta_{\ell}(R)\geqslant q^{n}+q^{n-1}+\cdots+1$. On the other hand, $\eta_{\ell}(R)\leqslant q^{n}+q^{n-1}+\cdots+1$ because the collection of all left ideals $L_{v}$ and $N_{V}$ forms a cover of $R$ by Lemmas 4.4(3) and 4.5(4). We conclude that $\eta_{\ell}(R)=q^{n}+q^{n-1}+\cdots+1=\frac{q^{n+1}-1}{q-1}$. ∎

### v2

@@ 4. An infinite family of $\eta_{\ell}$-elementary rings
For each $v\in\mathbb{F}_{q}^{n}$ (viewed as a column vector) and 1-dimensional subspace $V\subseteq\mathbb{F}_{q}^{n}$, let $L_{v}:=\{(A|Av):A\in S\}$ and $N_{V}:=\{(A|w):AV=0\text{ and }w\in\mathbb{F}_{q}^{n}\}$. That is, $N_{V}$ consists of the elements $(A|w)\in R$ such that $V$ is contained in the right nullspace of $A$, and there are no restrictions on $w$.
@@
    For each 1-dimensional subspace $V\subseteq\mathbb{F}_{q}^{n}$, $N_{V}$ is a maximal left ideal of $R$.
@@
    For each 1-dimensional subspace $V\subseteq\mathbb{F}_{q}^{n}$, $N_{V}$ is a cyclic left ideal of $R$.
@@
    Let $V$ and $W$ be 1-dimensional subspaces of $\mathbb{F}_{q}^{n}$. Then, $N_{V}=N_{W}$ if and only if $V=W$.
@@
    Let $(A|v)\in R$ such that $A\in S$ is not invertible. Then, $(A|v)\in N_{V}$ for any 1-dimensional subspace $V\subseteq\mathbb{F}_{q}^{n}$ such that $AV=0$.
@@
(1) Let $V$ be a 1-dimensional subspace of $\mathbb{F}_{q}^{n}$. It is easy to verify that $N_{V}$ is a left ideal of $R$. For maximality, let $V^{\perp}$ be the subspace complement to $V$ in $\mathbb{F}_{q}^{n}$. Then, $\dim_{\mathbb{F}_{q}}(V^{\perp})=n-1$, so $|V^{\perp}|=q^{n-1}$. Now, $(A|w)\in N_{V}$ if and only if each row of $A$ is the transpose of a vector in $V^{\perp}$. Thus, there are $q^{(n-1)n}$ choices for $A$ and $q^{n}$ choices for $w$, so $|N_{V}|=q^{(n-1)n}q^{n}=q^{n^{2}}$. Hence, $N_{V}$ is maximal by Lemma 4.2.
(2) Fix a 1-dimensional subspace $V\subseteq\mathbb{F}_{q}^{n}$ with complement subspace $V^{\perp}$. Let $\{b_{1},\ldots,b_{n-1}\}$ be a basis (of column vectors) for $V^{\perp}$. Let
@@
Then, $b_{i}^{T}v=0$ for all $1\leqslant i\leqslant n-1$ and all $v\in V$, so $B\in N_{V}$. Next, given $(A|w)\in N_{V}$, the transpose of row $i$ of $A$ is in $V^{\perp}$, so that row is an $\mathbb{F}_{q}$-linear combination of $b_{1}^{T},\ldots,b_{n-1}^{T}$. For $1\leqslant i\leqslant n$ and $1\leqslant j\leqslant n-1$, let $c_{i,j}\in\mathbb{F}_{q}$ be such that row $i$ of $A$ is equal $c_{i,1}b_{1}^{T}+\cdots+c_{i,n-1}b_{n-1}^{T}$. Also, let $w_{i}$ be the entry of $w$ in row $i$. Then,
@@
(3) This follows from the observation that for any 1-dimensional subspaces $V$ and $W$, there exist matrices $A_{V}$ and $A_{W}$ such that the right nullspaces of $A_{V}$ and $A_{W}$ are exactly $V$ and $W$, respectively.
@@
By Lemmas 4.4(1), 4.5(1) and 4.5(2), each left ideal $L_{v}$ and $N_{V}$ is a maximal left ideal of $R$, and is generated (as a left ideal) by a single element. Thus, each $L_{v}$ and $N_{V}$ must be part of any cover of $R$ by left ideals. By Lemma 4.4(2), the number of distinct maximal left ideals $L_{v}$ is $|\mathbb{F}_{q}^{n}|=q^{n}$; and by Lemma 4.5(3), the number of distinct maximal left ideals $N_{V}$ is equal to the number of 1-dimensional subspaces of $\mathbb{F}_{q}^{n}$, which is (using Gaussian binomial coefficients) $\binom{n}{1}_{q}=1+q+\cdots+q^{n-1}$. So, $\eta_{\ell}(R)\geqslant q^{n}+q^{n-1}+\cdots+1$. On the other hand, $\eta_{\ell}(R)\leqslant q^{n}+q^{n-1}+\cdots+1$ because the collection of all left ideals $L_{v}$ and $N_{V}$ forms a cover of $R$ by Lemmas 4.4(3) and 4.5(4). We conclude that $\eta_{\ell}(R)=q^{n}+q^{n-1}+\cdots+1=\frac{q^{n+1}-1}{q-1}$. ∎
@@ 5. Structure of noncommutative $\eta_{\ell}$-elementary rings
6. Covering numbers of modules
In this section, we prove Theorem 1.5, which determines all possible finite covering numbers of modules by submodules. Throughout, we work with left modules, although our proofs and techniques apply equally well to right modules.
Notation.
For a unital ring $R$ and a left $R$-module $M$, we let $\sigma_{R}(M)$ be the cardinality of a minimal cover of $M$ by proper left $R$-submodules.
Lemma 6.1.
Let $R$ be a unital ring and let $M$ be a left $R$-module that admits a finite cover by proper left $R$-submodules. Then, there exist a residue ring $\overline{R}$ of $R$ and a quotient module $\overline{M}$ of $M$ such that $\overline{R}$ is a finite semisimple ring, $\overline{M}$ is a left $\overline{R}$-module, and $\sigma_{R}(M)=\sigma_{\overline{R}}(\overline{M})$.
Proof.
Following the proof of [13, Thm. 2.2(2),(3)], we may assume that $M$ and $R$ are both finite. As in [13], we may assume that $M$ is finite by considering the induced cover of the underlying abelian group. Furthermore, we may view $M$ as a faithful left $R/{\rm{ann}}_{R}(M)$-module, but since $M$ is finite, this is a finite quotient. So, without loss of generality we may assume that $R$ is also finite.
Let $J=\mathscr{J}(R)$ be the Jacobson radical of $R$, let $\overline{R}=R/J$, and let $\overline{M}$ be the left $\overline{M}$-module $M/JM$. Since $M$ is finitely generated, $JM$ equals the intersection of all maximal $R$-submodules of $M$; and so any minimal irredundant cover of $M$ (by maximal $R$-submodules) can be projected to a minimal cover of $\overline{M}$ by $\overline{R}$-submodules. Conversely, a minimal cover of $\overline{M}$ can be lifted to a minimal cover of $M$. Thus, $\sigma_{R}(M)=\sigma_{\overline{R}}(\overline{M})$. ∎
Lemma 6.2.
Let $R$ be a finite semisimple ring and let $M$ be a left $R$-module that admits a finite cover by left $R$-submodules. Then, there exists a simple direct summand $S$ of $R$ and an $R$-submodule $N\subseteq M$ such that $N$ is a left semisimple $S$-module and $\sigma_{R}(M)=\sigma_{S}(N)$.
Proof.
Since $R$ is semisimple, we may write $R=\bigoplus_{i=1}^{t}S_{i}$ for some $t\geqslant 1$, where each $S_{i}$ is a finite simple ring. For each $i$, let $M_{i}=S_{i}M$, so that $M=\bigoplus_{i=1}^{t}M_{i}$. If $M_{i}$ is a cyclic $S_{i}$-module for each $i$, then $M$ is not coverable. So, at least one $M_{i}$ must admit a cover by proper $S_{i}$-submodules. We may now apply [13, Prop. 2.5] (which is the analog of Lemma 3.1 for modules) and conclude that there exists $1\leqslant j\leqslant t$ such that $\sigma_{R}(M)=\sigma_{S_{j}}(M_{j})$. ∎
Proof of Theorem 1.5.
By Lemmas 6.1 and 6.2, we may assume that $R$ is a finite simple ring and $M$ is a finite left $R$-module. So, $R=M_{n}(\mathbb{F}_{q})$ for some prime power $q$ and an integer $n\geqslant 1$. The $R$-module $M$ is semisimple, so we may write $M=\bigoplus_{i=1}^{\lambda}L_{i}$, where $\lambda\geqslant 1$ and each $L_{i}$ is a simple $R$-module. Now, each $L_{i}$ is isomorphic (as a left $R$-module) to $\mathbb{F}_{q}^{n}$. Treating the elements of each $L_{i}$ as column vectors, we may express $M$ as the set of all $n\times\lambda$ matrices with entries from $\mathbb{F}_{q}$, and treat the action of $R$ on $M$ as matrix multiplication.
If $1\leqslant\lambda\leqslant n$, then $M$ is the cyclic $R$-module generated by the $n\times\lambda$ matrix where each diagonal entry is 1 and all other entries are 0. Since $M$ is coverable, we must have $\lambda\geqslant n+1$. In this case, we may view $M$ as a nonunital ring with a left identity such that the left ideals of $M$ are precisely the left $R$-submodules of $M$. Indeed, embed $M$ in the matrix ring $M_{\lambda}(\mathbb{F}_{q})$ by sending each $m\in M$ to the $\lambda\times\lambda$ matrix $\binom{m}{0}$, where the final $\lambda-n$ rows are all zero. One may then check (cf. Lemma 4.1) that $L\subseteq M$ is a left ideal of $M$ if and only if $L$ is a left $R$-submodule of $M$. Thus, $\sigma_{R}(M)=\eta_{\ell}(M)$. Finally, the only $\eta_{\ell}$-elementary residue rings of $M$ are isomorphic to the ring of Theorem 1.3(1). Therefore, $\sigma_{R}(M)=\tfrac{q^{n+1}-1}{q-1}$. ∎
Acknowledgments
The authors thank the referee for their feedback and careful reading of the paper.
Disclosure Statement
The authors have no competing interests to declare.